### v1

# An analytical exploration of the optomechanical attractor diagram and of limit cycles

Abstract
We analyse the interplay between mechanical and radiation pressure in an optomechanical cavity system. Our study is based on an analytical evaluation of the infinite Bessel summations involved, which previously had led to a numerical exploration of the so-called attractor diagram. The analytical expressions are then suitable for further asymptotic analysis in opposing regimes of the amplitude, which allows for a characterisation of the diagram in terms of elementary functions. The same method is applied to the study of optomechanical limit cycles, extending previous results for enhanced values of the parameter space without the need for approximations such as the resolved sideband approximation.
I Introduction
Maxwell predicted that electromagnetic radiation can exert forces on material objects [1]. Typically, this force is negligible, but it has important effects in specific situations, such as in the tails of comets or during star formation. Optical forces have been used since the 1970s to trap and manipulate small particles and even individual atoms; see [2] for example, for early references. There is currently great interest in using radiation forces to manipulate the centre of mass motion of mechanical oscillators, ranging from the nanoscopic to the macroscopic (LIGO project) [3]. The phenomenon of damping can be described as a process whereby the motion of a system is slowed down by the application of a force. This force can be either positive or negative, with the former being responsible for cooling the motion and the latter for amplifying small forces [4].
While cooling motion is one of the cornerstones of cavity optomechanics, we focus here on the amplification features. Indeed, the light-induced forces can also result in a negative contribution to the overall damping rate. Initially, this increases the mechanical quality factor (Q) of the mechanical degree of freedom (for example, a cantilever), thereby amplifying the response to any noise source acting on the cantilever. Once the overall ‘damping rate’ becomes negative, which can occur simply by increasing the light power entering the cavity, the system no longer exhibits damping behaviour but instead displays instability [5, 6].
In the initial section, we present a methodology for the analytical study of the ensuing attractor diagram that has been numerically outlined in references [7, 8]. The determination of radiation pressure represents a fundamental problem in optomechanical systems and the force and power input are well-known to be characterised by infinite sums of Bessel functions [7, 3]. As we will see at the end of the Outlook section, these infinite summations are ubiquitous in a wide range of problems in cavity optomechanics and magnomechanics and beyond, including an extremely large number of quantum Floquet problems. This paper will exploit that these sums can be actually computed exactly [9, 10] and discuss a number of interesting predictions from the exact formulas, focusing exclusively on the dynamical multistability [7, 8, 11] and limit cycle problems in cavity optomechanics [12, 13].
Indeed, optomechanical systems are known to exhibit self-sustaining limit cycles in which the quantum state of the mechanical resonator exhibits non-classical properties [12]. The second part studies such limit cycles. We explain here tentatively some of the basic aspects of [12], that introduced several simplifications over the usual master equation [3, 14] describing the optomechanical system. In particular, an effective equation of motion of the Fokker-Planck type was derived in [12] for the motion of the mirror, based on the assumption that the cavity dynamics adiabatically follows the mechanical oscillator (assumption valid when the cavity decay rate $\kappa$ is larger than the characteristic coupling strength of the oscillator and the cavity mode). Our analytical findings apply to all cases presented in [12, 13]. It should be stressed that other works at the same time carried out similar analysis, for example [15, 16], using a truncated Wigner function.
One of the applications of the effective equation of motion is the subsequent description of optomechanical limit cycles in the quantum regime. The objective of this paper as it concerns to limit cycles is to characterise the drift and diffusion coefficients in [12] again through the utilisation of exact evaluations of Bessel summations, previously employed in [9, 10]. The results will be compared with those presented in [12], which did not consider the exact summation of the formulas. This led them to consider a drive setting close to resonance, specifically in the sideband-resolved regime with a detuning close to the resonance. This resonance implies that only the first terms in the sums need to be retained. We investigate the differences and further possibilities afforded by considering the summation in full, analytically. For example, some limit cycles that appear in [12] turn out to be an artifact of the approximation as they disappear as solutions in the exact formula. We will also specifically examine the implications of the exact summation of the diffusion coefficient for the Wigner function.
II Radiation pressure: parametric instability and Cooling
In the last decades the classical theory of the optomechanical parametric instability produced by radiation pressure backaction has been extensively studied in a large number of works, starting with the influential experimental works [17, 4]. We shall follow the very well-known ensuing theoretical analysis of [7, 11], where a numerical analysis of the analytical expressions of the model were carried out, culminating in a rather pervasive attractor diagram, that has also been portrayed in many reviews and works afterwards [18]. A consequence of our analytical formulas obtained below will be a quantitative analysis of this attractor diagram for certain regions of the parameter space. These results apply to rather generic optomechanical systems, consisting of a driven optical cavity and a movable mirror attached to a cantilever.
The analysis in [7] relied on a simple ansatz to determine the cavity’s amplitude $A$. In particular,
$x(t)=\bar{x}+A\cos(\Omega_{{\rm m}}t)\ ,$ (1)
for the self-induced oscillations. The radiation pressure force $F(t)$ is deduced by solving the classical equation for the light field amplitude, which is
$\dot{\alpha}=-\frac{\kappa}{2}(\alpha-\alpha_{{\rm max}})+i(\Delta+Gx(t))\alpha\,,$ (2)
where $\alpha_{{\rm max}}$ is defined to be the amplitude reached inside the cavity right at resonance, we have $\alpha_{{\rm max}}=2\alpha_{{\rm in}}\sqrt{\kappa_{{\rm ex}}}/\kappa$.
The time-averaged power input due to this force is $\left\langle F\dot{x}\right\rangle$. After inserting the ansatz (1), the solution for the optical degrees of freedom can be obtained as a Fourier series $\alpha(t)=e^{i\varphi(t)}\sum_{n}\alpha_{n}e^{in\Omega_{{\rm m}}t}$, with coefficients [7]
$\alpha_{n}=\frac{\alpha_{{\rm max}}}{2}\frac{J_{n}(-\frac{GA}{\Omega_{{\rm m}}})}{in\frac{\Omega_{{\rm m}}}{\kappa}+\frac{1}{2}-i(G\bar{x}+\Delta)/\kappa}\,,$ (3)
where the global phase is $\varphi(t)=(GA/\Omega_{{\rm m}})\sin(\Omega_{{\rm m}}t)$. Now the force $F(t)=\hbar G\left|\alpha(t)\right|^{2}$ and the time averages $\left\langle F\dot{x}\right\rangle$ and $\left\langle F\right\rangle$ can be calculated. One has [7, 3]
$\left\langle\left|\alpha(t)\right|^{2}\right\rangle=\sum_{n}\left|\alpha_{n}\right|^{2},$ (4)
and
$\left\langle\left|\alpha(t)\right|^{2}\dot{x}\right\rangle=A\Omega_{{\rm m}}{\rm Im}\sum_{n}\alpha_{n}^{*}\alpha_{n+1}.$ (5)
II.1 Analytical characterization
These series have always been dealt with numerically. Following the same analytical results employed while studying superconducting qubits [9] (see the Appendix for a summary), we can now give their full analytical form. For the time-averaged force we obtain
$\left\langle F(t)\right\rangle=\hbar G\left\langle\left|\alpha(t)\right|^{2}\right\rangle=\hbar G\frac{\kappa\big{|}\alpha_{{\rm max}}\big{|}^{2}}{2\Omega_{{\rm m}}}{\rm Im}\left(\frac{\pi}{\sin(\pi\rho)}\,J_{\rho}(X)J_{-\rho}(X)\right)\ ,$ (6)
where
$\qquad X\equiv-\frac{GA}{\Omega_{{\rm m}}}\ ,\ \ \rho\equiv\frac{1}{\Omega_{m}}(G\bar{x}+\Delta+\frac{i\kappa}{2})\ .$ (7)
To evaluate (5), we first write
$\alpha_{n}^{*}\alpha_{n+1}=\frac{\kappa^{2}\big{|}\alpha_{{\rm max}}\big{|}^{2}}{4\Omega_{{\rm m}}^{2}}\ \frac{J_{n}(X)J_{n+1}(X)}{(n-r+ic)(n+1-r-ic)}\ ,\qquad r\equiv\frac{1}{\Omega_{{\rm m}}}(G\bar{x}+\Delta)\ ,\ \ c\equiv\frac{\kappa}{2\Omega_{{\rm m}}}\ .$ (8)
Thus
$\alpha_{n}^{*}\alpha_{n+1}=\frac{\kappa^{2}\big{|}\alpha_{{\rm max}}\big{|}^{2}}{4\Omega_{{\rm m}}^{2}}\ \frac{J_{n}(X)J_{n+1}(X)}{1-2ic}\left(\frac{1}{n-r+ic}-\frac{1}{n+1-r-ic}\right)\ .$ (9)
We can now apply the exact summation formula. We find, for the time-averaged power input due to the radiation force
$\left\langle F\dot{x}\right\rangle=\left\langle\left|\alpha(t)\right|^{2}\dot{x}\right\rangle=-A\frac{\kappa^{2}\big{|}\alpha_{{\rm max}}\big{|}^{2}}{4\Omega_{{\rm m}}}{\rm Im}\left[\frac{1}{1-2ic}\left(\frac{\pi}{\sin(\pi\nu)}J_{\nu}(X)J_{1-\nu}(X)+\frac{\pi}{\sin(\pi\nu^{*})}J_{1+\nu^{*}}(X)J_{-\nu^{*}}(X)\right)\right].$ (10)
with $\nu\equiv-r+ic$.
Similarly as in the previous examples, one can give the asymptotic formulas at $|X|\gg 1$, $|X|\gg|\nu|^{2}$. In particular, we find for the force average:
$\left\langle F(t)\right\rangle=\hbar\frac{\kappa\big{|}\alpha_{{\rm max}}\big{|}^{2}}{A}b\left(\cosh(\frac{\pi\kappa}{2\Omega_{{\rm m}}})-\cos(\pi r)\sin(\frac{2GA}{\Omega_{{\rm m}}})\right)\ ,\qquad\frac{GA}{\Omega_{{\rm m}}}\gg 1,$ (11)
with
$b\equiv\frac{\sinh(\frac{\pi\kappa}{2\Omega_{{\rm m}}})}{\cosh(\frac{\pi\kappa}{\Omega_{{\rm m}}})-\cos(2\pi r)}\ ,$ (12)
and
$\left\langle F\dot{x}\right\rangle=\left\langle\left|\alpha(t)\right|^{2}\dot{x}\right\rangle=A\frac{\kappa^{2}\big{|}\alpha_{{\rm max}}\big{|}^{2}}{4\Omega_{{\rm m}}}\frac{8c\cosh(\pi c)\sin(\pi r)\cos(2X)}{\left(4c^{2}+1\right)X(\cos(2\pi r)-\cosh(2\pi c))}$ (13)
Thus the functional dependence of the time-averaged force satisfies properties for special values of the parameters. In particular, it is worth noticing the change of periodicity in $r$ occurring at $2GA=n\pi\Omega_{{\rm m}}$, $n\in\mathbb{Z}$, the suppression of oscillations in $A$ occurring at $2GA=(n+\frac{1}{2})\pi\Omega_{{\rm m}}$ and the expected resonances occurring at integer values of $r$ and small $c$.
II.2 Parametric instability, amplification and cooling: The attractor diagram
The condition of force balance determines the average position of the oscillator, thereby yielding an implicit equation for $\bar{x}$ [7],
$\langle\ddot{x}\rangle\equiv 0\quad\Leftrightarrow\quad m\omega_{M}^{2}\bar{x}=\left\langle F_{{\rm rad}}\right\rangle=\frac{\hbar g}{mx_{\text{ZPF}}}\langle\left|\alpha(t)\right|^{2}\rangle\,,$ (14)
where the average radiation force, $\langle F_{\text{rad}}\rangle$ is a function of the parameters $\bar{x}$ and $A$.
The balance between the mechanical power gain due to the light-induced force, $P_{{\rm rad}}=\left\langle F_{{\rm rad}}\dot{x}\right\rangle$, and the frictional loss $P_{{\rm fric}}=\Gamma_{M}\left\langle\dot{x}^{2}\right\rangle$ follows from [7, 11]
$\langle\dot{x}\ddot{x}\rangle\equiv 0\quad\Leftrightarrow\quad\langle F_{\text{rad}}\dot{x}\rangle=\Gamma_{M}\langle\dot{x}^{2}\rangle.$ (15)
For each value of the oscillation amplitude $A$ the plot of the ratio between radiation power input and friction loss, $P_{{\rm rad}}/P_{{\rm fric}}=\langle F_{\text{rad}}\dot{x}\rangle/(\Gamma_{M}\langle\dot{x}^{2}\rangle)$ is analyzed in [7, 11]. Therefore, the study of the ratio has been subjected to graphical and numerical analysis as has been also reviewed on numerous occasions after [7, 11]. The power balance is fulfilled if the ratio is one, whereas if the power input into the cantilever by radiation pressure is greater than frictional losses (i.e., for a ratio greater than one), the amplitude of oscillations will increase, otherwise it will decrease. Thus, stable solutions (dynamical attractors) are given by decrements in the ratio while increasing oscillation amplitude (energy).
It is now possible to study these features analytically in different regimes, for which simplified analytical expressions can be obtained by either asymptotic or power series expansions of the summed expressions. Experimental exploration of the optomechanical attractor diagram and its dynamics was carried out in [19].
II.2.1 Large driving amplitudes and resolved/unresolved sideband limit
The expression (13) describes the setting of large amplitudes. The region is the outer parabola given by $|X|\gg|\nu|^{2}$ in the $P_{{\rm rad}}=0$ plane of the attractor diagram. After the identification $X\equiv-\frac{GA}{\Omega_{{\rm m}}}$, (13) reads
$\left\langle F\dot{x}\right\rangle=\left\langle\left|\alpha(t)\right|^{2}\dot{x}\right\rangle=-\frac{\kappa^{2}\big{|}\alpha_{{\rm max}}\big{|}^{2}}{4G}\frac{8c\cosh(\pi c)\sin(\pi r)\cos(2X)}{\left(4c^{2}+1\right)(\cos(2\pi r)-\cosh(2\pi c))},$ (16)
which shows that the dependence with the amplitude $X$ is of zero order, whereas for the friction pressure is always quadratic. Therefore, for very large amplitudes mechanical friction pressure will always dominate and the ratio goes to 00, as the attractor diagram also clearly shows. Intermediate amplitudes are still well described by the asymptotic formula (16) if $\mu$ is not too large. Notice also that, in general, for very large amplitudes, the starting Ansatz itself (1) is not correct anyway [7].
In [7, 11], it is observed that for sufficiently strong driving, self-induced oscillations appear around integer multiples of the cantilever frequency, $\Delta\approx n\omega_{M}$. Notice that our solution is generically oscillatory both in the detuning $r$ and the amplitude $X$ (the only dependence on the amplitude) for all of its domain of validity. The denominator term $(\cos(2\pi r)-\cosh(2\pi c))$ is always negative hence the overall sign depends on the $\sin(\pi r)$ and $\cos(2X)$ terms.
An important parameter in optomechanical cavities is the sideband resolution ratio, defined as the ratio of the mechanical resonance frequency to the linewidth of the optical mode [11]. That is, $1/c$ where $c$ was given in (8). This parameter determines whether the optical mode effectively interacts with one or both sidebands of the mechanical mode. The former occurs for $c\sim 1$ and is typically referred to as the resolved sideband regime, while the latter occurs for an optical linewidth comparable to or greater than the mechanical resonance frequency and is referred to as the unresolved sideband regime (or bad cavity limit [20], $c\gg 1$). Sideband resolution is a measure of the dynamic interactions between the optical and mechanical degrees of freedom.
The attractor diagram in [7, 11] has been studied for the resolved sideband regime. This regime is of interest because it is crucial for ground-state cooling [8, 21], and $c=1/20$ has already been realised experimentally in [22]. For $r=(2m+1)/4$ and $m=0,1,2,...$ the two sideband resolution limits give (for simplicity we write the case $m=0$):
$\left\langle F\dot{x}\right\rangle$ $\simeq$ $\frac{\kappa^{2}\big{|}\alpha_{\mathrm{max}}\big{|}^{2}}{4G}\frac{\sqrt{2}e^{-\pi c}\cos(2X)}{c}\text{ for }c\gg 1,$
$\left\langle F\dot{x}\right\rangle$ $\simeq$ $\frac{\kappa^{2}\big{|}\alpha_{\mathrm{max}}\big{|}^{2}}{4G}4\sqrt{2}c\cos(2X)\text{ for }c\ll 1.$
The functional dependence with $c$ and $X$ is the same for other choices of $r$ that do not vanish the pressure (that is, other than $r$ integer). For example, for $r=1/2$, then:
$\left\langle F\dot{x}\right\rangle$ $\simeq$ $\frac{\kappa^{2}\big{|}\alpha_{\mathrm{max}}\big{|}^{2}}{4G}\frac{2e^{-\pi c}\cos(2X)}{c}\text{ for }c\gg 1,$
$\left\langle F\dot{x}\right\rangle$ $\simeq$ $\frac{\kappa^{2}\big{|}\alpha_{\mathrm{max}}\big{|}^{2}}{4G}4c\cos(2X)\text{ for }c\ll 1.$
II.2.2 Small driving amplitudes
A good approximation to steady-state properties of the system can be obtained for small driving amplitudes $X<<1$, normally valid for weak optomechanical coupling. In this way, we explore another region of the attractor diagram [7, 11], namely the one near the origin of the energy and detuning plane. We first can show that the radiation pressure is quadratic in the amplitude $X$. Since it is also quadratic in the amplitude for the friction this leads to a comparison of the two corresponding coefficients.
It follows immediately by having the Bessel function evaluation. Using [Watson, 5.41, Eq1] we have:
$J_{\nu}(x)J_{\mu}(x)=\sum_{m=0}^{\infty}\frac{(-1)^{m}\left(\frac{x}{2}\right)^{\mu+\nu+2m}\left(\mu+\nu+m+1\right)_{m}}{m!\Gamma\left(\nu+m+1\right)\Gamma\left(\mu+m+1\right)}.$
At first order, it is linear in $x$ and that makes the radiation pressure quadratic in the amplitude, for small amplitudes as expected. We have at first order:
$\frac{\pi}{\sin\pi\nu}J_{\nu}(x)J_{1-\nu}(x)$$\approx$ $\frac{x}{2\nu\left(1-\nu\right)},$
$\frac{\pi}{\sin\pi\nu^{\ast}}J_{1+\nu^{\ast}}(x)J_{-\nu^{\ast}}(x)$$\approx$ $\frac{x}{2\nu^{\ast}\left(1+\nu^{\ast}\right)}.$
Using these together with the expression obtained for the radiation pressure, recalling also that $\nu\equiv-r+ic$, we then have
$\left\langle F\dot{x}\right\rangle$$\approx$ $-A\frac{\kappa^{2}\big{|}\alpha_{\mathrm{max}}\big{|}^{2}}{4\Omega_{\mathrm{m}}}\mathrm{Im}\left[\frac{1}{1-2ic}\left(\frac{X}{2\nu\left(1-\nu\right)}+\frac{X}{2\nu^{\ast}\left(1+\nu^{\ast}\right)}\right)\right]$
$=$ $-GA^{2}\frac{\kappa^{2}\big{|}\alpha_{\mathrm{max}}\big{|}^{2}}{2\Omega_{\mathrm{m}}^{2}}\frac{cr}{\left(r^{2}+c^{2}\right)(c^{2}+(r-1)^{2})(c^{2}+(r+1)^{2})}$
Therefore, for stability, after dividing by the friction radiation $P_{fric}=\pi A^{2}\Omega_{\mathrm{m}}$, the stability criteria is that
$f(r)\equiv\frac{\left\langle F\dot{x}\right\rangle}{P_{fric}}=-2G\frac{\big{|}\alpha_{\mathrm{max}}\big{|}^{2}}{\pi\Omega_{\mathrm{m}}}\frac{c^{3}\,r}{\left(r^{2}+c^{2}\right)(c^{2}+(r-1)^{2})(c^{2}+(r+1)^{2})}<1\ .$ (17)
Cooling occurs when the ratio is negative since then the radiation pressure is negative. Recall that $c>0$ then the negative-pressure condition is simply
$r<0$
which implies the simple condition
$\frac{1}{\Omega_{\mathrm{m}}}\left(G\overline{x}+\Delta\right)<0\ .$
Let us now consider the implications of the bound (17). A close examination of the function $f(r)$ shows that the maximum value occurs for $c=1/\sqrt{2}$ at $r=c$, where
$f(x_{\rm max})\bigg{|}_{c_{\rm max}}=\frac{G\big{|}\alpha_{\mathrm{max}}\big{|}^{2}}{4\pi\Omega_{\mathrm{m}}}<1$ (18)
This condition therefore ensures stability for all values of $c$. On the other hand, for a given $c$ the stability condition on the parameter $G\big{|}\alpha_{\mathrm{max}}\big{|}^{2}/\Omega_{\mathrm{m}}$ may be weaker.
Figure 1: The maximum value of the ratio $f(r)\equiv\frac{\left\langle F\dot{x}\right\rangle}{P_{fric}}$ as a function of $c$ (in units of $2G\big{|}\alpha_{\mathrm{max}}\big{|}^{2}/\pi\Omega_{\mathrm{m}}$).
For example, for small values of $c$ the absolute maximum is located at $r^{2}_{0}\approx 1-2c^{2}$. At the maximum, the bound becomes
$f(r_{0})\bigg{|}_{c\ll 1}=\frac{G\big{|}\alpha_{\mathrm{max}}\big{|}^{2}}{2\pi\Omega_{\mathrm{m}}}\ c<1$ (19)
This reproduces the small $c$ behavior of fig. 1.
For $c\gg 1$, the function $f(r)$ has the maximum located at $r_{1}\approx c^{2}/5+O(1)$, and the stability condition becomes
$f(r_{1})\bigg{|}_{c\gg 1}=\frac{25\sqrt{2}}{108}\,\frac{G\big{|}\alpha_{\mathrm{max}}\big{|}^{2}}{\pi\Omega_{\mathrm{m}}}\ \frac{1}{c^{2}}<1\ ,$ (20)
in agreement with fig. 1. As $c\gg 1$, this implies a weaker condition on the parameters $G\big{|}\alpha_{\mathrm{max}}\big{|}^{2}/\Omega_{\mathrm{m}}$. Indeed we have checked that the condition (19) arising at small values of $c$ is sufficient to ensure stability at all values of $c$ (at least in the small amplitude regime assumed in this subsection).
The equation obtained implies that for blue-detuning $\Delta>0$ one could also have cooling is the mechanical off-set $\overline{x}$ is such that $G\overline{x}<-\Omega_{\mathrm{m}}\Delta$. If we discuss, as in [11], in terms purely of the detuning $\Delta$ with the off-set $\overline{x}=0$. Then the condition is simply that red-detuning  [^1] For red-detuning $\Delta<0$ one can typically have cooling. To be more precise, the theory of resonator dynamics indicates that in this regime the energy tends to be transferred from the resonator to the cavity [7]. The back-action damping is positive, guaranteeing that the system will remain in a state where the mechanical oscillator fluctuates about a fixed point. leads to cooling whereas blue-detuning to amplification, as usual.
III Limit cycles
The same mathematical considerations can be applied to the study of limit cycles in [12]. An effective equation of motion, in the form of a Fokker-Planck (FP) equation was derived using a general parametrized phase-space quasiprobability distribution, which can be reduced to the P representation, Q (Husimi) representation, or the Wigner function $W$. The FP equation in [12] is transformed into polar coordinates, and the focus is on the time evolution of the oscillator amplitude $r$ by integrating out the phase variable $\phi$  [^2] The drift and diffusion coefficients of the Fokker-Planck equation are independent of the phase of the mechanical phase-space variable $\beta$ due to the rotating-wave approximation used in [12].. This results in a one-dimensional Fokker-Planck equation for the $Q$-function of the mechanical oscillator
$\dot{Q}(r)=-\partial_{r}\mu(r)Q(r)+\partial_{r}^{2}D(r)Q(r)\ ,$ (21)
where $\mu(r)$ and $D(r)$ are the drift and diffusion coefficients. It is worth mentioning that, while in the main text in [12] the case of the $Q$-function was considered, the more general parametrized phase-space quasiprobability distribution was presented as well and the drift coefficient mantains the same form regardless of the initial quasiprobability distribution used to describe the system’s density operator. However, the diffusion coefficient is specific to the phase-space distribution [12], but it is equally amenable to our Bessel summation formulas, as we show below for the case of the Wigner function.
III.1 Drift and diffusion coefficients of the effective equation of motion
We start by providing an exact calculation of the infinite sums describing the drift and diffusion coefficients. They are given by [12]
$\mu(r)$ $=$ $-\gamma r+g_{0}E^{2}\sum_{n}{\rm Im}\left[\frac{J_{n-1}(\eta r)J_{n}(\eta r)}{h_{n-1}h_{n}^{*}}\right]\ ,$ (22)
$D(r)$ $=$ $\frac{1}{2}\gamma(\bar{n}+1)+\frac{g_{0}^{2}E^{2}}{2}\sum_{n}\left(\frac{\kappa J_{n}(\eta r)^{2}}{|h_{n}|^{2}|\tilde{h}_{n-1}|^{2}}-{\rm Re}\left[\frac{J_{n-2}(\eta r)J_{n}(\eta r)}{\tilde{h}_{n-1}h^{*}_{n-2}h_{n}}\right]\right)\ ,$ (23)
where
$h_{n}=\kappa+i(n\omega_{m}-\Delta_{\rm eff})\ ,\qquad\tilde{h}_{n}=\kappa+i(n\omega_{m}-\tilde{\Delta}_{\rm eff})\ .$ (24)
Here $r$ represents the amplitude, $E=\hbar\omega_{m}r^{2}$ is the energy, $g_{0}$ is the single-photon coupling strength and the dimensionless coupling $\eta=2g_{0}/\omega_{m}$ is also used to characterize the optomechanical coupling [12], $\kappa$ is the cavity amplitude decay rate and $\Delta=\omega_{L}-\omega_{c}$ is the detuning between the laser frequency $\omega_{L}$ and the bare cavity resonance $\omega_{c}$. The mechanical damping rate parameter is denoted by $\gamma$.
The term effective detuning $\Delta_{\rm eff}$ represents the fact that there is a dynamical dependence of the effective detuning on the intracavity amplitude and a corresponding nonlinear dependence on the cycle amplitude. In particular:
$\Delta_{eff}=\Delta+2K\sum_{n}\left|\alpha_{n}\right|^{2},$
where $K$ denotes the strength of an effective Kerr nonlinearity with $K=\frac{g_{0}^{2}}{\omega_{m}}$. Indeed, in the formalism of [12] the cavity acquires an effective Kerr nonlinearity that is encoded in $\Delta_{eff}$. Because the Kerr nonlinearity induces a different effective detuning for the mean field $\alpha$ than for the fluctuations, we also have:
$\tilde{\Delta}_{eff}=\Delta+4K\sum_{n}\left|\alpha_{n}\right|^{2}.$
The static approximation for the detuning is valid for a negligible Kerr parameter $K$ or in the weak driving limit. Indeed, in the $\left|\alpha(t)\right|^{2}\ll 1$ limit the bare detuning is just shifted by the constant Kerr term and there is no need to distinguish between $\Delta_{eff}$ and $\widetilde{\Delta}_{eff}$. The shift $\Delta_{eff}=\Delta+K$ appears naturally in the framework of [12] (see also [16, 25]) and explains the appearance of limit cycles for specific particular values given in [12] and also, for example, for limit cycles on the blue sideband in [25]. So, already in the static approximation model, this classical Kerr shift is of importance.
A discussion of both the static and the dynamical setting is alternatively presented throughout the whole discussion in the paper [12]. Here, we first discuss the static case, valid for $\left|\alpha(t)\right|^{2}\ll 1$ and then in Section B.1 below, we discuss the dynamical case by evaluating $\Delta_{\rm eff}(r)$ in full, in contrast to the sideband-resolved approximation of [12]. Indeed, these sums have been evaluated in a close-to-resonance approximation [12], where you discard an infinite number of terms. Using the summation formulas of appendix A, we can now provide exact expressions for the complete resummation. For the drift, we obtain
$\mu(r)=-\gamma r-\frac{\pi g_{0}E^{2}}{\omega_{m}}{\rm Re}\left[\frac{1}{2\kappa-i\omega_{m}}\left(\frac{J_{\nu^{*}}(\eta r)J_{1-\nu^{*}}(\eta r)}{\sin(\pi\nu^{*})}+\frac{J_{1+\nu}(\eta r)J_{-\nu}(\eta r)}{\sin(\pi\nu)}\right)\right]\ ,$ (25)
with
$\nu\equiv\frac{\Delta_{\rm eff}}{\omega_{m}}+\frac{i\kappa}{\omega_{m}}\ .$ (26)
Details of the calculation are given in appendix A and below we will discuss its asymptotic value for large $\eta r$. For the diffusion coefficient, we find
$D(r)=\frac{1}{2}\gamma(\bar{n}+1)+D_{a}(r)+D_{b}(r)\ ,$ (27)
where
$D_{a}(r)=-\frac{\pi}{2\omega_{m}}\,g_{0}^{2}E^{2}{\rm Im}\left[B\left(\frac{J_{\nu}(\eta r)J_{-\nu}(\eta r)}{\sin(\pi\nu)}+\frac{J_{1+\tilde{\nu}^{*}}(\eta r)J_{-1-\tilde{\nu}^{*}}(\eta r)}{\sin(\pi\tilde{\nu}^{*})}\right)\right]$ (28)
$D_{b}(r)=\frac{\pi}{2\omega_{m}}\,g_{0}^{2}E^{2}{\rm Im}\left[C_{1}\frac{J_{\nu}(\eta r)J_{2-\nu}(\eta r)}{\sin(\pi\nu)}+C_{2}\frac{J_{1+\tilde{\nu}^{*}}(\eta r)J_{1-\tilde{\nu}^{*}}(\eta r)}{\sin(\pi\tilde{\nu}^{*})}+C_{3}\frac{J_{2+\nu^{*}}(\eta r)J_{-\nu^{*}}(\eta r)}{\sin(\pi\nu^{*})}\right]$ (29)
$\tilde{\nu}=\frac{\tilde{\Delta}_{\rm eff}}{\omega_{m}}+\frac{i\kappa}{\omega_{m}}$ and the constants $B$, $C_{1}$, $C_{2}$, $C_{3}$ are given in appendix A.
III.1.1 Diffusion in the Wigner case
The Fokker-Planck equation (21) can also be written for the Wigner function. It has the same drift coefficient, and a diffusion given by (see (B.20) in [12])
$D_{W}=\frac{1}{4}\gamma(2\bar{n}+1)+\frac{1}{4}g_{0}^{2}E^{2}\sum_{n}\frac{\kappa}{|\tilde{h}_{n+1}|^{2}}\left(\frac{J_{n+2}J_{n+2}}{h_{n+2}h_{n+2}^{*}}+\frac{J_{n}J_{n}}{h_{n}h_{n}^{*}}-\frac{J_{n}J_{n+2}}{h_{n}h_{n+2}^{*}}-\frac{J_{n}J_{n+2}}{h_{n}^{*}h_{n+2}}\right).$ (30)
All the arguments of the Bessel functions are $\eta r$. This expression plays a relevant role in the dynamics of limit cycles. The infinite sums can again be computed exactly by using the summation formulas of the appendix. The final result is a long expression and we will not need it here, but $D_{W}$ drastically simplifies when $\tilde{\Delta}_{\rm eff}=\Delta_{\rm eff}$. This identification is justified in the sideband-resolved regime studied in [12]. We obtain (see Appendix for details)
$D_{W}=\frac{1}{4}\gamma(2\bar{n}+1)+\frac{1}{4}\,g_{0}^{2}E^{2}\left(D_{W}^{(1)}+D_{W}^{(2)}+D_{W}^{(3)}+D_{W}^{(3)*}\right)\ ,$ (31)
where
$D_{W}^{(1)}=-\frac{\pi}{\omega_{m}^{2}}{\rm Im}\left[\frac{J_{-\nu}J_{\nu}}{\beta\,\sin(\pi\nu)}-\frac{J_{\nu+1}J_{-\nu-1}}{\beta^{*}\,\sin(\pi\nu)}\right]\ ,\qquad\beta=\omega_{m}-2i\kappa\ ,$
$D_{W}^{(2)}=-\frac{\pi}{\omega_{m}^{2}}{\rm Im}\left[\frac{J_{-\nu}J_{\nu}}{\beta^{*}\,\sin(\pi\nu)}-\frac{J_{\nu-1}J_{-\nu+1}}{\beta\,\sin(\pi\nu)}\right]\ ,$
$D_{W}^{(3)}=\frac{\pi}{2\omega_{m}^{2}(\omega_{m}+2i\kappa)}\left[\frac{\kappa}{\omega_{m}+i\kappa}\left(\frac{J_{\nu+2}J_{-\nu}}{\sin(\pi\nu)}-\frac{J_{\nu^{*}}J_{-\nu^{*}+2}}{\sin(\pi\nu^{*})}\right)-\frac{iJ_{\nu+1}J_{-\nu+1}}{\sin(\pi\nu)}+\frac{iJ_{\nu^{*}+1}J_{-\nu^{*}+1}}{\sin(\pi\nu^{*})}\right]\ ,$
where $\nu$ has been defined in (26).
These exact formulas may be compared with the approximation in Equation (48) of [12], valid close to the resonance point $n=0$. This approximation requires $\Delta_{\rm eff}\ll\omega_{m}$ in which case the original sum (30) reduces to the $n=0$ term,
$D_{W}^{(0)}=\frac{1}{4}\gamma(2\bar{n}+1)+\frac{\kappa g_{0}^{2}E^{2}}{\omega_{m}^{4}}\left(J_{1}^{2}(\eta r)+\frac{\omega_{m}^{2}}{2(\kappa^{2}+\Delta_{\rm eff}^{2})}\,J_{0}^{2}(\eta r)\right)\ .$ (32)
In fig 2 we compare our exact formula for $D_{W}$ with $D_{W}^{(0)}$ for $\kappa=0.1$, $\omega_{m}$ (resolved sideband) and two different values of $\Delta_{\rm eff}$. We see that the approximation (32) is very inaccurate for $\Delta_{\rm eff}=0.6\,\omega_{m}$ A closer examination shows that the approximation (32) is no longer applicable (with an error higher than 20 per cent) for $\Delta_{\rm eff}>0.4\,\omega_{m}$.
On the other hand, the exact representation (31) holds for any value of the parameters, and therefore can be applied even for very high values of $\Delta_{\rm eff}$ and $\kappa$. For large values of $r$, one can use the asymptotic formulas for the Bessel functions and express the full result in terms of trigonometric functions. Examples of this procedure are given in the next section.
Refer to caption   Refer to caption
(a) (b)
Figure 2: Optical part of the Wigner diffusion from approximation (32) (blue) and for the exact formula (31) (red) for $\kappa=0.1\omega_{m}$. a) $\Delta_{\rm eff}=0.1\,\omega_{m}$. The approximation (32) reproduces with high accuracy the exact curve. b) $\Delta_{\rm eff}=0.6\,\omega_{m}$. The approximation (32) is no longer applicable as higher $n$ terms become important.
III.2 Asymptotic formulas, classical optomechanics cycles and the dynamical setting $\Delta\equiv\Delta_{\rm eff}(r)$
In order to see the implications of these formulas, it is useful to look at the asymptotic behavior for large arguments of the Bessel function, in a parameter regime where the full resummation is important. That is, for frequencies away from resonance: in this regime all secondary resonances significantly contribute to the sum. For large $x$, the Bessel functions have the familiar asymptotic expression [26]
$J_{\nu}(x)\approx\frac{\sqrt{2}}{\sqrt{\pi x}}\,\cos(x-\nu\frac{\pi}{2}-\frac{\pi}{4})\ .$ (33)
Using this formula, we can find a revealing expression for the drift (25) in the regime of large $\eta r$. We obtain
$\mu(r)=-\gamma r-\frac{8E^{2}g_{0}\kappa\sin\left(\frac{\pi\Delta}{\omega_{m}}\right)\cosh\left(\frac{\pi\kappa}{\omega_{m}}\right)\cos(2\eta r)}{\eta r\omega_{m}\left(4\kappa^{2}+\omega_{m}^{2}\right)\left(\cosh\left(\frac{2\pi\kappa}{\omega_{m}}\right)-\cos\left(\frac{2\pi\Delta}{\omega_{m}}\right)\right)}\ .$ (34)
To simplify the notation, here we wrote $\Delta\equiv\Delta_{\rm eff}(r)$. This formula exhibits the relevant periods of oscillations in $r$ and in $\Delta/\omega_{m}$. A surprising effect predicted by the asymptotic formula for $\mu(r)$ is that the second term in (34) identically vanish for
$\Delta=n\omega_{m},\qquad n\in\mathbb{Z}\ .$ (35)
Thus, for these special values of $\Delta$, $\mu(r)=-\gamma r+O(1/r)$. Oscillations are suppressed by a factor $1/r$ originating from subleading terms in the asymptotic expansion of the Bessel functions. This effect is shown in fig. 3 and is a result of the full resummation. It is not seen if one truncates the sum to few values of $n$, as in a close-to-resonance approximation. It would be interesting to see this effect in experiments.
Refer to caption Refer to caption Refer to caption
(a) (b) (c)
Figure 3: The exact $\mu(r)$ for two different values of $\Delta$. Oscillations are suppressed by an extra factor $1/r$ when $\Delta=n\omega_{m}$. Here $\gamma=0$, $\kappa=\omega_{m}=1$. (a) In blue, $\Delta=0.8\omega_{m}$, in red $\Delta=\omega_{m}$. (b) Plot of $r\mu(r)$ for $\Delta=0.8\omega_{m}$, exhibiting the $\mu\sim 1/r$ asymptotic behavior when $\Delta/\omega_{m}$ is not an integer. (c) Plot of $r^{2}\mu(r)$ exhibiting the $\mu\sim 1/r^{2}$ behavior for $\Delta=\omega_{m}$.
III.2.1 Classical optomechanical limit cycles and a comparison with close to resonance results
The theory for classical optomechanical limit cycles, as presented in [12], is described by just the drift component (25) of the Fokker-Planck equation. In the absence of diffusion and with a constant effective detuning $\Delta_{\rm eff}$, the evolution of the oscillator amplitude is deterministic, and the oscillator amplitude satisfies the following equation [12]
$\dot{r}=\mu(r)=-r\ \gamma_{\rm eff}(r)\ ,\qquad\gamma_{\rm eff}(r)=\gamma+\gamma_{\rm opt}(r)\ .$ (36)
The combined intrinsic and optically induced damping of the oscillator is denoted by $\gamma_{\rm eff}(r)$. The amplitudes $r_{0}$ for limit cycles are determined by the conditions $\gamma_{\rm eff}(r_{0})=0$ and $\gamma_{\rm eff}^{\prime}(r_{0})>0$. Taking into account the complete sum defining $\mu(r)$ derived above in (25), we get an explicit formula for $\gamma_{\rm opt}(r)$ which therefore allows us to accurately determine the limit cycles.
This can be compared with the approximation used in (43) of [12] for the oscillator close to resonance,
$\gamma_{\rm opt}^{(0)}(r)=-\frac{g_{0}^{2}E^{2}}{\omega_{m}^{2}}\ \frac{2\kappa\Delta_{\rm eff}}{\Delta_{\rm eff}^{2}+\kappa^{2}}\,\frac{1}{r}\,J_{0}(\eta r)J_{1}(\eta r)\ .$ (37)
Refer to caption   Refer to caption
(a) (b)
Figure 4: Effective damping $\gamma_{\rm eff}$ as a function of the amplitude in units of $\gamma_{0}=g_{0}E^{2}/\omega_{m}^{2}$. Limit cycles corresponds to the zeros. (a) $\gamma=0$. The approximation (37) (brown) compared with the $\gamma_{\rm eff}$ obtained from the complete sum (blue). We see that the location of the zeros is significantly corrected when all the terms of the sum are included. (b) $\gamma=0.01\gamma_{0}$. In this case the approximated $\gamma$ (in brown) exhibits three limit cycles occurring at $r\approx 2.4$, $r\approx 5.3$ and $r\approx 8.1$, but the last limit cycle disappears in the more accurate description using (25) (in blue).
In contrast, $\gamma_{\rm opt}(r)$ with $\mu$ given by (25) keeps the contributions of all values of $n$, not only $n=0$ and $n=1$. The resulting plot has a similar shape. However, for some values of the parameters there are crucial differences, since some limit cycles that appear in the approximation (37) turn out to be an artifact of the approximation. Indeed, they disappear as solutions in the exact formula. This is shown in fig. 4, done for $\Delta_{\rm eff}=\kappa$, $\kappa=0.6\,\omega_{m}$. (cf. with fig. 1a) of [12]). Although for these values of parameters there are significant modifications in the limit cycles, in a sideband-resolved regime with $\Delta_{\rm eff},\ \kappa\ll\omega_{m}$ the differences obviously become smaller. Nonetheless, this example shows that the complete sum could be useful in a more accurate description of the limit cycles in a much wider range of parameters.
III.2.2 Dynamical setting
We can similarly make use of the summation formula to compute a formula for $\Delta_{\rm eff}(r)$ itself, in the dynamical setting. This is determined by a nonlinear equation [12]
$\Delta_{\rm eff}=\Delta+2KE^{2}\sum_{n}\frac{J_{n}(r\eta)J_{n}(r\eta)}{|\kappa+i(n\omega_{m}-\Delta_{\rm eff})|^{2}}\ .$ (38)
The infinite sum can again be computed using the summation formulas of the appendix by noting that this can be re-writen as
$\Delta_{\rm eff}=\Delta+\frac{iKE^{2}}{\kappa\omega_{m}}\sum_{n}\left(\frac{J_{n}(r\eta)J_{n}(r\eta)}{n-\nu}-\frac{J_{n}(r\eta)J_{n}(r\eta)}{n-\nu^{*}}\right)\ .$ (39)
with $\nu=\frac{\Delta_{\rm eff}}{\omega_{m}}+i\frac{\kappa}{\omega_{m}}$, as before (26). Now the summation formula gives
$\Delta_{\rm eff}=\Delta+\frac{2\pi KE^{2}}{\kappa\omega_{m}}{\rm Im}\left[\frac{J_{\nu}(r\eta)J_{-\nu}(r\eta)}{\sin(\pi\nu)}\right]\ .$ (40)
One can similarly study the asymptotic behavior at large $r$. In this regime $\Delta_{\rm eff}$ becomes
$\Delta_{\rm eff}\approx\Delta-\frac{4KE^{2}}{\kappa\omega_{m}\eta r}\,\frac{\sinh(\pi\frac{\kappa}{\omega_{m}})}{\cosh(2\pi\frac{\kappa}{\omega_{m}})-\cos(2\pi\frac{\Delta_{\rm eff}}{\omega_{m}})}\left(\cosh(\pi\frac{\kappa}{\omega_{m}})+\cos(\pi\frac{\Delta_{\rm eff}}{\omega_{m}})\sin(2r\eta)\right)\ .$ (41)
This can be solved by assuming that $\Delta_{\rm eff}\ll\omega_{m}$. Then
$\Delta_{\rm eff}\approx\Delta-\frac{2KE^{2}}{\kappa\omega_{m}\eta r}\,\frac{1}{\sinh(\pi\frac{\kappa}{\omega_{m}})}\left(\cosh(\pi\frac{\kappa}{\omega_{m}})+\sin(2r\eta)\right)\ .$ (42)
This formula encodes the information of all secondary resonances.
IV Outlook
We have seen that the radiation pressure of a monochromatic light source can be fully analysed. This extends the analytical understanding of such a central physical quantity [3, 7, 8] and allows a number of consequent considerations, both at the exact level and for asymptotic analysis for regions of parameter space that would otherwise be inaccessible. We have seen how the physics can then be analysed for the whole parameter space without the usual limitations which are due to the lack of evaluation of all the sideband contributions.
Indeed, the type of expressions we have evaluated, which in addition to the radiation pressure includes also the drift and diffusion coefficients, including Wigner diffusion, in the Fokker-Planck formulation of the effective equations of motion [12], are known to be quite ubiquitous in cavity optomechanics and in other problems of the Floquet type [9, 10]. Therefore, many other uses of the evaluations here can be devised, both for work along the lines we have considered here, such as [13, 16, 15], and for other settings and/or quantities not explored here but manifestly amenable to a similar or even analogous treatment. The set of examples includes the consideration of periodic modulation of optical resonators [27, 28] and the study of scattering of N-photons in a periodically modulated cavity [29], optomechanical systems with the inclusion of thermal effects [30], the generation of frequency combs both at the optic and at the magnonic level [30, 28, 31, 32, 33], the study of Floquet-Weyl semimetals [34], the use of analogies between the coupled sideband dynamics of optomechanical oscillations and a Bloch-band structure [35] or the study of photon-photon correlation function [36]. Recently, there has been considerable interest and progress in extending quantum amplification features by considering a Floquet setting [37, 38, 39]. The analytical evaluations presented here (and in related studies [9, 10]) can also be applied to further study spin Floquet amplification.
It is evident that the implementation of the summation formula, along with its exact properties and asymptotics, would prove to be of significant benefit to each of these research lines. On the other hand, the feasibility of applying these methods to other problems, involving synchronization or to the multiple drive setting, such as those outlined in [40, 41] for example, appears to be a distinct possibility in principle but it is important to note that the analytical study for example of equilibrium conditions in [40] or the generalized Bessel setting of [41], is more arduous.
The technical advance of this paper lies in the formulation of a compact, simple formula that comprehensively encompasses all terms within the summation. This provides an analytical understanding of the parameter space and offers insight into the underlying physics of the various processes. Additionally, it facilitates the application of asymptotic expressions, resulting in expressions that are expressed in terms of uniquely elementary functions. Since these are simple trigonometric functions, further elucidation, for example of the zeros [10], becomes a particularly straightforward task, helping to identify special points in regions of the parameter space in limits where a resolved sideband approximation, which retains only a few resonant terms in the summation, does not reach. There are also open problems already in the context of the present discussion. For example, whether the more complete analysis of Wigner diffusion presented here also allows a more detailed analysis of energy fluctuations by evaluating the Fano factor, as in [12], but with full summation. All these considerations require no further mathematical developments than those presented here or in the mathematically related works [9, 10].
A novel extension, such as the consideration of multichromatic instead of monochromatic light, would imply the substitution of Bessel functions by generalised Bessel functions [42, 43]. Progress along these lines is possible and will be reported elsewhere, even though the fundamental summation formula does not extend naively to this setting. The interest in this extended setting goes beyond optomechanics, magnomechanics or optical resonators applications as it is also a fundamental extension in a number of quantum Floquet problems with periodic modulations [44, 27, 41], including the consideration of a number of quantum technologies platforms where having multiple drives is essential for the control of the quantum system.
Acknowledgements.
We thank Professors David Pérez-García, Oriol Romero-Isart and Patrick Maurer for discussions and correspondence at different stages of this work. The work of MT was started at Universidad Complutense de Madrid (UCM) and continued and finished after moving to the Shanghai Institute for Mathematics and Interdisciplinary Sciences (SIMIS). JGR acknowledges financial support from grant 2021-SGR-249 (Generalitat de Catalunya) and by the Spanish MCIN/AEI/10.13039/501100011033 grant PID2022-126224NB-C21. MT acknowledges previous financial support from FEI-EU-22-06, funded by Universidad Complutense de Madrid, and grant PID2020-113523GB-I00, funded by the Spanish Ministry of Science and Innovation.
Appendix A Bessel summations and diffusion coefficients
In order to compute the infinite sums in the drift and diffusion coefficient expressions in (23), we will need the following summation formulas [9, 45]:
$\sum_{n}\frac{J_{n-1}(x)J_{n}(x)}{n+\mu}$ $=$ $-\frac{\pi}{\sin(\pi\mu)}\,J_{-\mu}(x)J_{\mu+1}(x)$
$\sum_{n}\frac{J_{n}(x)J_{n}(x)}{n+\mu}$ $=$ $\frac{\pi}{\sin(\pi\mu)}\,J_{-\mu}(x)J_{\mu}(x)$
$\sum_{n}\frac{J_{n-2}(x)J_{n}(x)}{n+\mu}$ $=$ $\frac{\pi}{\sin(\pi\mu)}\,J_{-\mu}(x)J_{\mu+2}(x)$ (43)
We first write the denominator in (22) as follows
$\frac{1}{h_{n-1}h^{*}_{n}}=\frac{1}{2\kappa-i\omega_{m}}\left(\frac{1}{h_{n-1}}+\frac{1}{h^{*}_{n}}\right)\ .$ (44)
Substituting this expression into (22), a direct application of the summation formulas (43) yields the compact analytic expression (25).
A.1 Diffusion coefficient
Now, in addition to the formulas above, it is convenient to write the expression as sum of terms of with simpler denominators linear in $n$. We have
$\frac{1}{|h_{n}|^{2}|\tilde{h}_{n-1}|^{2}}=\frac{B}{2\kappa}\left(\frac{1}{h_{n}}+\frac{1}{\tilde{h}^{*}_{n-1}}\right)+\frac{B^{*}}{2\kappa}\left(\frac{1}{h_{n}^{*}}+\frac{1}{\tilde{h}_{n-1}}\right)$ (45)
$\frac{1}{h_{n}{\tilde{h}}_{n-1}h^{*}_{n-2}}=\frac{C_{1}}{h_{n}}-\frac{C_{2}}{\tilde{h}_{n-1}}-\frac{C_{3}}{h_{n-2}^{*}}$ (46)
with
$B\equiv\frac{1}{(\omega_{m}+\tilde{\Delta}_{\rm eff}-\Delta_{\rm eff})(\omega_{m}+\tilde{\Delta}_{\rm eff}-\Delta_{\rm eff}-2i\kappa)}$ (47)
$C_{1}\equiv\frac{1}{2(\omega_{m}+\tilde{\Delta}_{\rm eff}-\Delta_{\rm eff})(\omega_{m}-i\kappa)}\ ,\quad C_{2}\equiv\frac{1}{(\omega_{m}+\tilde{\Delta}_{\rm eff}-\Delta_{\rm eff})(\omega_{m}-\tilde{\Delta}_{\rm eff}+\Delta_{\rm eff}-2i\kappa)}\ ,$ (48)
$C_{3}\equiv\frac{1}{2(\omega_{m}-i\kappa)(\omega_{m}-\tilde{\Delta}_{\rm eff}+\Delta_{\rm eff}-2i\kappa)}$ (49)
Now the formulas (43) can be directly applied. In the particular case $\tilde{\Delta}_{\rm eff}=\Delta$, this gives the closed expression (31).
A.2 Wigner diffusion
Once again the sums can be computed exactly using the summation formulas (43). As in the previous cases, the strategy is to write each fraction as sum of individual terms containing $1/h_{n}$ or $1/\tilde{h}_{n}$. For example, for the first term,
$\frac{\kappa}{|\tilde{h}_{n+1}|^{2}|h_{n+2}|^{2}}=\frac{1}{2(\omega-\Delta+\tilde{\Delta})}\left(\frac{A}{h_{n+2}}+\frac{A^{*}}{h_{n+2}^{*}}+\frac{A^{*}}{\tilde{h}_{n+1}}+\frac{A}{\tilde{h}_{n+1}^{*}}\right)\ ,$ (50)
with
$A\equiv\frac{1}{\omega-\Delta+\tilde{\Delta}-2i\kappa}\ ,$
and similarly for the rest of the terms.